OpenStudy (anonymous):

Complete the square and write the equation in standard form. Then give the center and radius of the circle. x2 + y2 - 10x - 8y + 29 = 0 A. (x - 5)2 +(y - 4)2 = 12 (5, 4), r = 12 B. (x - 5)2 +(y - 4)2 = 12 (-5, -4), r = 2 C. (x - 5)2 +(y - 4)2 = 12 (5, 4), r = 2 D. (x + 5)2 +(y + 4)2 = 12 (-5, -4), r = 2

OpenStudy (cwrw238):

(x - 5)^3 - 25 + (y - 4)^2 - 16 + 29 = 0 (x - 5)^3 + (y - 4)^2 - 41 + 29 = 0 (x - 5)^3 + (y - 4)^2 = 12 thats A

OpenStudy (anonymous):

thnx :)

OpenStudy (cwrw238):

yw